# Unrolled Creative Adversarial Network For Generating Novel Musical Pieces

### v1

@@ Abstract
Music generation has been established as a prominent topic in artificial intelligence and machine learning [2, 13] over recent years. In most recent works [4, 7, 3], RNN-based neural network methods have been applied for sequence generation. In contrast, generative adversarial networks (GANs) and their counterparts have been explored by very few researchers [10] for music generation.
In this paper, a classical system was employed alongside a new system to generate creative music. Both systems were designed based on adversarial networks to generate music by learning from examples. The classical system was trained to learn a set of music pieces without differentiating between classes, whereas the new system was trained to learn the different composers and their styles to generate a creative music piece by deviating from the learned composers’ styles.
The base structure utilized was generative adversarial networks (GANs), which are capable of generating novel outputs given a set of inputs to learn from and mimic their distribution. It has been shown in previous work that GANs are limited in their original design with respect to creative outputs [5]. Building on the Creative Adversarial Networks (CAN) [5], this work applied them in the music domain rather than the visual art domain. Additionally, unrolled CAN was introduced to prevent mode collapse. Experiments were conducted on both GAN and CAN for generating music, and their capabilities were measured in terms of deviation from the input set.
@@ 1 Introduction
Machines have always been shown as devices capable of replicating certain learned behaviours. It has been a dream to make creative machines and is of much desire in the machine learning scope and artificial intelligence. After all if we want machines to mimic human behaviour, one the most important human traits is creativity. Take driving for example; humans are capable of using their creativity and creative problem solving skills, to avoid unexpected road events, preventing accidents. That is a skill that all self driving car developers hope to accomplish. The art of generating music is a very complex problem [10] and is a different challenge than that of text generation. It includes finding a rhythm between the musical notes and learning polyphony and many more challenges.
To study style in music lets take a look at what style means in language. When one speaks or writes their style is defined by their selection of words an vocabulary, their choice of grammar and linguistics, all governed within constraints of the language and dialect one has learned, and have not created. However, since constraints can allow for interpretations of different realizations it produces variety. Therefore, a pattern in a style of writing or a piece of music may not relate in every aspect, only some aspects are selected to convey pattern replication. 1,2 shows two examples of musical pieces replicating a pattern of using pitches (8-7/2-1) typically preferred and chosen by classical musicians [9].
Figure 1: Symphony No.46 in B major Haydn
@@
There are two clefs in a music sheets (3) that are used to differentiate between high and bass notes and are called the treble clef and the bass clef. The treble clef is usually played with the right hand on the piano and the bass clef is usually played with the left hand on the piano.
@@
In music sheets some note may be represented different graphically yet representing the same note. And sometimes certain music sheets may look identical graphically while actually are representing different notes. That is due to the change in the clef (Check figure 4).
A professional musician must be able to identify the similarities and differences in the notes. That comes naturally since professional musicians are regularly reading sheet music.
Figure 4: Same-Different exercise example [12]
@@ 2.2 Data Preprocessing
The data used in this project contain midi files. In order to pass them to the GAN or the CAN the midi files were converted to images by converting each note to a pixel location. The resulting image file can be converted back to midi file without losing any of the information. (see figure 5)
@@ 2.3 Music GAN architecture
A detailed architecture of Discriminator and Generator. 7,8
@@ 2.4 Music CAN and unrolled CAN architecture
In the music CAN the architecture[11] [5]is still made up of two neural networks that compete against each other. However, the loss here stresses on the generator motivating it to produce a musical piece that cannot be classified as belonging to one class. In other words, the discriminator should be capable of classifying the original real data to their composers or class label. The discriminator should not be able to find a suitable class for the generated piece of music. Thus the input data must be labeled by the class it belongs to in order to reduce the loss on real data classification. The generator architecture is similar to that of the GAN where it takes a random z vector as an input and generates an image representing the midi file.
@@
A detailed architecture of Discriminator and Generator. 10, 11
@@
3 Technical Details
3.1 Generative Adversarial Networks
GAN was first present by Ian J. Goodfellow on 2014 where two deep neural networks play a minmax game over the value function $V(G,D)$ [6]. Where G is the Generator and D is the Discriminator. And z is a random vector supplied to the generator as input.
[…]
$\mathbb{E}_{z\sim p_{z}(z)}[log(1-D(G(z)))]$ (1)
3.2 Creative Adversarial Networks
CAN was presented recently where a modification to the value function to include cross entropy loss over the classes of the generated output ($G(z)$)[5]. Thus now the architecture includes classification and therefore, the loss also includes classification loss. And in order to convey creativity the authors used class ambiguity loss. The class ambiguity loss is achieved by maximizing the class posterior entropy. And hence, the value function is as follows
[…]
$\mathbb{E}_{x,\hat{c}\sim p_{data}}[logD_{r}(x)+logD_{c}(c=\hat{c}|x)]+$
$\mathbb{E}_{z\sim p_{z}}[log(1-D_{r}(G(z)))-\sum_{k=1}^{K}(\frac{1}{K}log(D_{c}(c_{k}|G(z)))+$
$(1-\frac{1}{K})log(1-D_{c}(c_{k}|G(z))))]$
Where as above z here is the random vector. And c here represents the class label.
3.3 Unrolling Creative Adversarial Networks
The unrolled CAN is a combination between unrolled GAN and CAN architecture where this structure allows the generator network to look into the future of the discriminator network and optimize over the future discriminator but not updating the discriminator just yet[8]. In other words, the generator is a number (unrolling step numbers may vary) of steps ahead than the discriminator. The result is a reduction in the strength of the discriminator allowing the generator to catch up. Figure 12 shows the result of the regular CAN network after 20 epochs and the results after the modification was introduced making it an unrolled CAN.
Figure 12: Example of CAN, with k=0, output converging to almost identical result after 20 epochs
The way the parameters are updated in GAN is described by the set of equations below
$G\leftarrow G-\eta\frac{dV(D(G),G)}{dG}$ (3)
$D\leftarrow D-\eta\frac{dV(D,G)}{dD}$ (4)
where $V(D,G)$ is the value function. The updates in parameters after unrolling will be as follows
$G\leftarrow G-\eta\frac{dV(D^{k}(G),G)}{dG}$ (5)
$D\leftarrow D-\eta\frac{dV(D,G)}{dD}$ (6)
Where $\eta$ here represents the learning rate. and $k$ is the number of unrolling steps. See figure 13 for the changes due to Unrolled-CAN.
Figure 13: Example of unrolled CAN, with k=1, output converging to varying result after 20 epochs
@@ 4.1 Scoring novelty generated musical pieces
In this paper, the expert gate technique [1] was followed, where an auto-encoder was trained on the reconstruction of the real dataset. The expert gate was trained to become proficient in reconstructing the real dataset, which was then used to evaluate the novelty of the generated images. The Mean Squared Error (MSE) was used as the scoring metric. Higher scores indicated greater creativity and deviation from the training set. After the training process, the reconstruction error of the real dataset was observed to have an average of 0.01 MSE. The network architecture is shown in Figure 14.
@@ 4.5 Discussion
Table 1: Results. Unrolled CAN is better.
Non the less the CAN architecture was very good at generating music that deviates from the composers’ styles but converging to the generation of similar results leading to mode collapse which could be the reason why the average reconstruction error was very low.
4.6 Reproducing the Results
@@
To train the Unrolled CAN model on the dataset, the file GAN_reconstructed.ipynb located in the folder CAN_models should be executed. Sample data has been provided in the classical_data_midi folder. The files training.py and model.py contain all the model-related configurations. MIDI datasets can be preprocessed into images using the data_pre_processing.ipynb file.
@@
The novelty_score.ipynb file has been utilized to calculate the novelty score, as discussed in [1]. This file is located in the home directory.
@@ 5 Conclusion
This paper explored different ways musical styles are classified and reviewed multiple architectures. It was demonstrated that the GAN structure [6] and CAN [5] are fully capable of generating music. Additionally, a novel modification to the CAN was introduced by unrolling it, as detailed in [8]. This project combined art and science, leveraging machine learning to create novel and innovative music. Such advancements could inspire new playlists for hotels, aid singers and music composers in seeking new inspirations, and reaffirm the purpose of machines as tools to benefit humanity.

### v3

# https://pratiknag.com/

@@ Abstract
Music generation has emerged as a significant topic in artificial intelligence and machine learning. While recurrent neural networks (RNNs) have been widely employed for sequence generation, generative adversarial networks (GANs) remain relatively underexplored in this domain. This paper presents two systems based on adversarial networks for music generation. The first system learns a set of music pieces without differentiating between styles, while the second system focuses on learning and deviating from specific composers’ styles to create innovative music. By extending the Creative Adversarial Networks (CAN) framework to the music domain, this work introduces unrolled CAN to address mode collapse, evaluating both GAN and CAN in terms of creativity and variation.
@@ 1 Introduction
Machines have long been envisioned as tools capable of replicating learned behaviors. However, instilling creativity in machines remains a formidable challenge in the fields of machine learning and artificial intelligence. Creativity, a quintessential human trait, is pivotal in tasks such as problem-solving and decision-making. For instance, in self-driving car development, enabling machines to make creative decisions is essential for handling unforeseen road events and preventing accidents.
Music generation is particularly complex, involving challenges like learning polyph- ony, rhythm, and intricate patterns of musical notes [muhamed2021symbolic]. Style in music, akin to style in language, is governed by constraints. In language, style emerges from vocabulary selection, grammar, and linguistic preferences. Similarly, in music, patterns of style manifest through pitch choices, rhythms, and dynamics, producing variety within certain constraints [meyer1996style]. For instance, Figures 1 and 2 illustrate patterns in classical music that replicate specific stylistic elements preferred by renowned composers.
Figure 1: Symphony No.46 in B major, Haydn
@@
Sheet music also presents challenges in interpretation. Notes represented graphically may appear different but sound the same, or conversely, identical-looking sheets may represent different notes due to clef changes [schon2001naming]. Professional musicians rely on experience to identify such nuances naturally. Figures 3 and 4 demonstrate these complexities, including examples of treble and bass clefs and exercises that highlight the "same-different" note identification challenge.
@@
Figure 4: Same-Different exercise example [schon2001naming]
Recent advancements have explored RNN-based approaches for music generation, focusing on sequence prediction tasks [yang2017midinet, briot2020deep]. On the other hand, research into GAN-based frameworks has been comparatively limited [huang2020pop, davis2014generating, chen2001creating]. Building on these developments, this work adapts the Creative Adversarial Networks (CAN) framework [DBLP] from visual arts to the music domain. By introducing unrolled CAN, this study addresses challenges like mode collapse and evaluates the potential of GAN and CAN in creative music generation.
2 Technical Details
2.1 Generative Adversarial Networks
Generative Adversarial Networks (GANs), first introduced by Goodfellow et al. (2014) [goodfellow2014generative], consist of two neural networks engaged in a two-player minimax game: a Generator $G$ and a Discriminator $D$. The generator maps a noise vector
$z\sim p_{z}(z),$
typically drawn from a simple prior distribution (e.g., Gaussian or uniform), to the data space to produce synthetic samples $G(z)$. The discriminator receives either real data samples
$x\sim p_{\text{data}}(x)$
or generated samples $G(z)$, and outputs a scalar value $D(x)\in[0,1]$ representing the probability that the input is real.
The minimax value function for the original GAN is defined as
$\min_{G}\max_{D}V(D,G)=\mathbb{E}_{x\sim p_{\text{data}}(x)}[\log D(x)]+\mathbb{E}_{z\sim p_{z}(z)}[\log(1-D(G(z)))].$ (1)
The discriminator attempts to maximize this objective by improving its ability to distinguish real from generated samples, while the generator attempts to minimize it by producing samples that the discriminator classifies as real.
2.2 Creative Adversarial Networks
Creative Adversarial Networks (CANs) extend the GAN formulation by incorporating stylistic classification and explicit class ambiguity objectives [DBLP]. In this setup, the discriminator outputs two components:
  • $D_{r}(x)$: the probability that $x$ is real,
  • $D_{c}(c\mid x)$: the posterior distribution over stylistic classes $c\in\{1,\dots,K\}$.
For real samples, the discriminator is trained both to classify them as real and to correctly identify their style class $\hat{c}$. For generated samples $G(z)$, the goal is to enforce class ambiguity by maximizing the entropy of the predicted class posterior. This prevents the generator from reproducing known styles and encourages creativity. The value function for CAN is therefore:
$\begin{split}\min_{G}\max_{D}V(D,G)=\;&\mathbb{E}_{x,\,\hat{c}\sim p_{\text{data}}}\!\left[\log D_{r}(x)+\log D_{c}(\hat{c}\mid x)\right]\\ &+\mathbb{E}_{z\sim p_{z}(z)}\Bigg[\log\!\left(1-D_{r}(G(z))\right)\\ &\qquad-\sum_{k=1}^{K}\!\left(\frac{1}{K}\log D_{c}(c_{k}\mid G(z))+\left(1-\frac{1}{K}\right)\!\log(1-D_{c}(c_{k}\mid G(z)))\right)\Bigg].\end{split}$ (2)
The last term enforces high-entropy outputs over the class distribution, ensuring that the generated samples do not collapse into any known stylistic class.
2.3 Unrolling Creative Adversarial Networks
Unrolling Creative Adversarial Networks (Unrolled-CAN) combine the CAN architecture with the unrolled optimization procedure introduced in Unrolled GANs [metz2017unrolled]. The core idea is to allow the generator to anticipate how the discriminator will change in future training steps, thereby smoothing the adversarial dynamics and reducing instances of mode collapse. In a standard GAN or CAN, parameters are updated as:
$G\leftarrow G-\eta\frac{\partial V(D,G)}{\partial G},$ (3)
$D\leftarrow D-\eta\frac{\partial V(D,G)}{\partial D},$ (4)
where $\eta$ denotes the learning rate.
In an unrolled model, the discriminator is virtually updated for $k$ unrolling steps (without committing these updates to the actual discriminator). The generator then optimizes the loss with respect to this $k$-step unrolled discriminator, denoted $D^{(k)}$. Thus, the updates become:
$G\leftarrow G-\eta\frac{\partial V(D^{(k)},G)}{\partial G},$ (5)
$D\leftarrow D-\eta\frac{\partial V(D,G)}{\partial D}.$ (6)
Here, $k$ is the number of unrolling steps. Increasing $k$ allows the generator to account for the discriminator’s future reaction, effectively weakening overly strong discriminators and enabling the generator to explore more diverse solutions.
Figures 5 and 6 illustrate the difference between a standard CAN ($k=0$) and an Unrolled-CAN ($k=1$), showing improved diversity and reduced convergence to a single mode after 20 epochs.
Figure 5: Example of standard CAN output with $k=0$, showing convergence to nearly identical samples after 20 epochs.
Figure 6: Example of Unrolled-CAN output with $k=1$, showing increased variability and reduced mode collapse after 20 epochs.
@@ 3.2 Data Preprocessing
The data used in this project contain midi files. In order to pass them to the GAN or the CAN the midi files were converted to images by converting each note to a pixel location. The resulting image file can be converted back to midi file without losing any of the information (see Figure 7).
@@ 3.3 Music GAN Architecture
A detailed architecture of Discriminator and Generator. 9,10
@@ 3.4 Music CAN and Unrolled CAN Architecture
In the music CAN the architecture [sbai2018design, DBLP] is still made up of two neural networks that compete against each other. However, the loss here stresses on the generator motivating it to produce a musical piece that cannot be classified as belonging to one class. In other words, the discriminator should be capable of classifying the original real data to their composers or class label. The discriminator should not be able to find a suitable class for the generated piece of music. Thus the input data must be labeled by the class it belongs to in order to reduce the loss on real data classification. The generator architecture is similar to that of the GAN where it takes a random z vector as an input and generates an image representing the midi file.
@@
A detailed architecture of Discriminator and Generator is shown in Figure 12 and 13.
@@ 4.1 Scoring Novelty Generated Musical Pieces
In this paper, the expert gate technique [aljundi2017expert] was followed, where an auto-encoder was trained on the reconstruction of the real dataset. The expert gate was trained to become proficient in reconstructing the real dataset, which was then used to evaluate the novelty of the generated images. The Mean Squared Error (MSE) was used as the scoring metric. Higher scores indicated greater creativity and deviation from the training set. After the training process, the reconstruction error of the real dataset was observed to have an average of 0.01 MSE. The network architecture is shown in Figure 14.
@@ 4.5 Discussion
Table 1: Results for different comparing models.
None the less the CAN architecture was very good at generating music that deviates from the composers’ styles but converging to the generation of similar results leading to mode collapse which could be the reason why the average reconstruction error was very low.
4.6 Reproducing Results
@@
To train the Unrolled CAN model on the dataset, the file GAN_reconstructed.- ipynb located in the folder CAN_models should be executed. The files training.py and model.py contain all the model-related configurations. MIDI datasets can be preprocessed into images using the data_pre_processing.ipynb file from the home directory.
@@
The novelty_score.ipynb file, placed in the home directory, has been utilized to calculate the novelty score, as discussed in [aljundi2017expert]. This file is located in the home directory.
@@ 5 Conclusion
This paper explored different ways musical styles are classified and reviewed multiple architectures. It was demonstrated that the GAN structure [goodfellow2014generative] and CAN [DBLP] are fully capable of generating music. Additionally, a novel modification to the CAN was introduced by unrolling it, as detailed in [metz2017unrolled]. This project combined art and science, leveraging machine learning to create novel and innovative music. Such advancements could inspire new playlists for hotels, aid singers and music composers in seeking new inspirations, and reaffirm the purpose of machines as tools to benefit humanity.